X3-2x is it odd function or even


1
Expert's answer
2021-05-07T10:21:43-0400

Let "f(x)=x^3-2x"

Require to verify whether the given function "f(x)" is even or odd.


Recollect the following:

A function "f(x)" is said to be even function if "f(-x)=f(x)" and

A function "f(x)" is said to be odd function if "f(-x)=-f(x)"


Now "f(x)=x^3-2x\Rightarrow f(-x)=(-x)^3-2(-x)"

"\Rightarrow f(-x)=-x^3+2x"

"\Rightarrow f(-x)=-(x^3-2x)"

"\Rightarrow f(-x)=-f(x)"

Therefore, the given function is odd function.